A 6.00 kg aluminum metal is suspended in a rope and partially immersed in water (πœŒπ‘€π‘Žπ‘‘π‘’π‘Ÿ = 1000 kg/m3). If the tension in the rope is 4.55 N, what volume of aluminum was immersed?Β 


1
Expert's answer
2022-06-06T08:55:24-0400
"F_b=mg-T=6.00*9.8-4.55=54.3\:\rm N"

"F_b=g\rho V_s"

"V_s=\frac{F_b}{g\rho}=\frac{54.3}{9.8*1000}=5.54*10^{-3}\;\rm m^3"